# Degree Distribution based Spiking Graph Networks for Domain Adaptation

### v3

@@ Appendix C Proof of Theorem 3
where (a) is the outcome of applying the union bound with coefficient $\frac{N^{\prime}_{T}}{N_{S}+N^{\prime}_{T}}$, $\frac{N_{S}}{N_{S}+N^{\prime}_{T}}$ respectively; (b) additionally adopt the assumption $N^{\prime}_{T}\ll N_{S}$, following the sleight-of-hand in Li et al. [2021a] Theorem 3.2.
@@
$+C\sqrt{\frac{ln(2/\delta)}{N_{S}}}+2C_{f}C_{g}W_{1}(\mathbb{P}_{S}(G),\mathbb{P}_{T}(G))+\omega^{\prime},$
where $\omega^{\prime}=\min_{||g||_{Lip}\leq C_{g},||f||_{Lip}\leq C_{f}}\{\epsilon_{S}(h,\hat{h})+\epsilon_{T}(h,\hat{h})\}$, $\hat{\epsilon}_{T}$ is the empirical risk on the high confidence samples, $\epsilon_{T}$ is the empirical risk on the target domain. Besides, we have:
@@
$+2C_{f}C_{g}W_{1}\left(\mathbb{P}_{S}(G),\mathbb{P}_{T}(G)\right)+\omega^{\prime}.$
Appendix D Algorithm
Algorithm 1 Learning Algorithm of DeSGDA
@@ Appendix D Algorithm
Output: Node degree-aware personalized spiking graph encoder parameters $\theta$, domain discriminator $\gamma$
1:Initialize model parameters $\theta$ and $\gamma$
@@
4:     Forward propagate $\mathcal{B}^{s}$ and $\mathcal{B}^{t}$ through the node degree-aware personalized spiking graph encoder
@@
7:     Update model parameters $\theta$ and $\gamma$ via backpropagation
@@
Appendix E Complexity Analysis
Here we analyze the computational complexity of the proposed DeSGDA. The computational complexity primarily relies on node degree-aware personalized spiking representations. For a given graph $G$, $\|A\|_{0}$ denotes the number of nonzeros in the adjacency matrix of $G$. $d$ is the feature dimension. $L$ denote the layer number of spiking graph encoder. $|V|$ is the number of nodes. $T$ denotes the number of time latency. The spiking graph encoder takes $\mathcal{O}\left(T\cdot L\cdot\left(\|A\|_{0}\cdot d+|V|\cdot d^{2}\right)\right)$ computational time for each graph. As a result, the complexity of our DeSGDA is proportional to both $|V|$ and $\|A\|_{0}$.
Figure 6: Visualization of different distributions on PROTEINS.
Figure 7: Visualization of different distributions on NCI1.
Figure 8: Visualization of different distributions on FRANKENSTEIN.
Figure 9: Visualization of different distributions on MUTAGNENICITY.
Appendix F Dataset
@@
PROTEINS 1,113 39.1 72.8 2
NCI1 4,110 29.87 32.30 2
MUTAGENICITY 4,337 30.32 30.77 2
FRANKENSTEIN 4,337 16.9 17.88 2
F.1 Dataset Description
@@
We conduct extensive experiments on four public benchmark graph datasets from TUDataset. The dataset statistics can be found in Table 3, and their details are shown as follows:
  • PROTEINS. The PROTEINS dataset Dobson and Doig [2003] consists of protein graphs. Each label indicates whether the protein graph is an enzyme. Nodes within these graphs represent amino acids connected by edges if neighbors along the amino acid sequence are spaced less than 6 Angstroms apart. We divide the PROTEINS dataset into four parts, i.e., P0, P1, P2, and P3, based on edge density, node density, and graph flux; the sub-datasets exhibit substantial domain disparities among them.
  • NCI1. The NCI1 Wale et al. [2008] dataset contains 4,100 chemical compounds with atoms as nodes and bonds as edges. Each label indicates the characteristics that hinder the growth of cancer cells. Like the PROTEINS dataset, we divide the NCI1 dataset into four parts, i.e., N0, N1, N2, and N3, based on edge density, node density, and graph flux.
  • FRANKENSTEIN. The FRANKENSTEIN Orsini et al. [2015] dataset consists of 4,337 molecular graphs, where nodes represent atoms and edges depict chemical bonds. Each graph is labeled to classify molecules based on their biological activity. Like the PROTEINS dataset, the entire FRANKENSTEIN dataset is divided into four segments (F0, F1, F2, and F3) based on edge density, node density, and graph flux.
  • MUTAGENICITY. The MUTAGENICITY Kazius et al. [2005] dataset contains 4,337 chemical compounds, each represented as a graph where nodes represent atoms and edges indicate bonds. Each graph can be used to identify mutagenic compounds, aiding studies in toxicology and chemical safety. Like the PROTEINS dataset, the entire MUTAGENICITY dataset is divided into four segments (M0, M1, M2, and M3) based on edge density, node density, and graph flux.
Table 4: The results of ablation studies on NCI1 (source → target). Bold results indicate the best performance per column.
@@
DeSGDA w/o CA 58.9 62.6 61.7 60.1 58.2 62.0 63.4 61.7 63.5 60.7 68.3 64.0 62.1
DeSGDA w/o PL 62.8 65.1 55.0 56.9 55.3 54.4 58.6 62.9 61.2 62.2 59.7 64.3 59.8
DeSGDA w/o CF 52.7 62.6 58.5 56.5 56.3 63.8 63.1 52.7 65.4 65.1 58.2 63.7 59.9
DeSGDA w/o TL 64.6 62.8 60.6 54.9 44.9 65.4 60.8 64.8 56.1 64.1 57.8 54.0 59.3
DeSGDA w PT 64.4 59.2 64.6 62.2 64.5 60.1 61.1 60.8 62.0 64.2 62.7 65.5 62.6
DeSGDA w CL 65.3 70.1 68.9 65.2 65.6 63.0 67.6 64.6 68.9 66.8 68.6 70.7 67.1
DeSGDA 68.5 71.4 70.1 69.0 68.9 66.3 69.6 70.2 71.1 69.3 74.4 70.0 69.9
Table 5: The results of ablation studies on FRANKENSTEIN (source → target). Bold results indicate the best performance.
@@
DeSGDA w/o CA 59.6 60.7 60.2 60.7 77.9 58.5 61.3 60.5 76.8 60.0 80.1 63.9 65.0
DeSGDA w/o PL 58.9 56.6 55.9 56.6 69.8 53.6 55.0 57.7 70.4 53.2 68.8 61.7 59.9
DeSGDA w/o CF 55.9 59.2 59.5 54.5 74.9 52.9 59.9 54.6 78.8 56.1 74.5 58.5 61.6
DeSGDA w/o TL 56.2 57.2 60.7 48.2 53.3 45.8 59.5 53.0 48.8 54.8 51.1 54.7 53.6
DeSGDA w PT 60.9 60.2 61.0 61.5 76.2 60.9 60.3 62.2 75.3 57.2 74.8 63.3 64.5
DeSGDA w CL 62.2 61.3 61.7 60.7 79.3 59.7 63.6 64.1 79.5 61.1 78.6 62.7 66.2
DeSGDA 63.5 64.1 62.6 63.0 81.1 62.4 65.5 65.4 81.9 60.8 82.0 65.9 68.2
Table 6: The results of ablation studies on MUTAGENICITY (source → target). Bold results indicate the best performance.
@@
DeSGDA w/o CA 58.3 63.3 62.1 60.1 81.4 62.7 63.8 59.1 80.3 57.1 78.9 63.0 65.8
DeSGDA w/o PL 58.8 57.6 60.2 61.4 73.3 54.5 63.7 61.0 61.2 61.9 71.0 60.2 62.1
DeSGDA w/o CF 54.7 58.3 59.6 58.7 72.1 60.7 57.9 54.9 78.0 57.4 63.2 60.4 61.3
DeSGDA w/o TL 56.5 60.7 59.1 60.7 43.4 56.3 51.0 56.7 53.3 56.9 63.1 52.5 55.9
DeSGDA w PT 62.1 64.3 63.9 61.8 77.3 62.4 66.6 63.9 73.5 66.1 82.1 64.9 67.4
DeSGDA w CL 66.1 62.6 63.3 64.7 81.5 60.7 69.8 67.0 82.2 67.8 82.4 63.6 69.3
DeSGDA 65.4 65.9 65.5 65.6 82.8 63.6 70.7 68.2 82.9 67.6 83.9 66.5 70.8
F.2 Data processing
In our implementation, we process these four TUDatasets by adding a self-loop connection for each node. Additionally, we utilize one-hot embeddings to represent node attributes for datasets (FRANKENSTEIN and MUTAGENICITY) where node features are unavailable.
Appendix G Baselines
Graph kernel method. We compare DeSGDA with one graph kernel method:
  • WL subtree: Weisfeiler-Lehman (WL) subtree Shervashidze et al. [2011] is a graph kernel method, which calculates the graph similarity by a kernel function, where It encodes local neighborhood structures into subtree patterns, efficiently capturing the topology information contained in graphs.
Graph neural networks. We compare DeSGDA with four widely used general graph neural networks:
  • GCN: GCN Kipf and Welling [2017] is a spectral-based neural network that iteratively updates node representations by aggregating information from neighboring nodes, effectively capturing both local graph structure and node features.
  • GIN: GIN Xu et al. [2018] is a message-passing neural network designed to distinguish graph structures using an injective aggregation function, theoretically achieving the expressive power of the Weisfeiler-Lehman test.
  • CIN: CIN Bodnar et al. [2021] extends the Weisfeiler-Lehman framework by integrating cellular complexes into graph neural networks, allowing for the capture of higher-dimensional topological features.
  • GMT: GMT Baek et al. [2021] utilizes self-attention mechanisms to dynamically adjust the importance of nodes based on their structural dependencies, thereby enhancing both adaptability and performance.
Spiking graph neural networks. We compare DeSGDA with two spiking graph neural networks:
  • SpikeGCN: SpikeGCN Zhu et al. [2022] introduces an end-to-end framework designed to integrate the fidelity characteristics of SNNs with graph node representations.
  • DRSGNN: DRSGNN Zhao et al. [2024] dynamically adapts to evolving graph structures and relationships through a novel architecture that updates node representations in real-time.
Domain adaption methods. We compare DeSGDA with two recent domain adaption methods:
  • CDAN: CDAN Long et al. [2018] employs a conditional adversarial learning strategy to reduce domain discrepancy by conditioning adversarial adaptation on discriminative information from multiple domains.
  • ToAlign: ToAlign Wei et al. [2021b] uses token-level alignment strategies within Transformer architectures to enhance cross-lingual transfer, optimizing the alignment of semantic representations across languages.
  • MetaAlign: MetaAlign Wei et al. [2021a] is a meta-learning framework for domain adaptation that dynamically aligns feature distributions across domains by learning domain-invariant representations.
Graph domain adaptation methods. We compare DeSGDA with six SOTA graph domain adaption methods:
  • DEAL: DEAL Yin et al. [2022b] uses domain adversarial learning to align graph representations across different domains without labeled data, overcoming discrepancies between the source and target domains.
  • CoCo: CoCo Yin et al. [2023a] leverages contrastive learning to align graph representations between source and target domains, enhancing domain adaptation by promoting intra-domain cohesion and inter-domain separation in an unsupervised manner.
  • SGDA: SGDA Qiao et al. [2023] utilizes labeled data from the source domain along with a limited amount of labeled data from the target domain to learn domain-invariant graph representations.
  • DGDA: DGDA Cai et al. [2024] employs generative models to capture the underlying distribution of graph data across domains, facilitating the transfer of graph structures and features by learning shared latent spaces.
  • A2GNN: A2GNN Liu et al. [2024a] introduces a novel propagation mechanism to enhance feature transferability across domains, improving the alignment of graph structures and node features in an unsupervised setting.
  • PA-BOTH: PA-BOTH Liu et al. [2024b] aligns node pairs between source and target graphs, optimizing feature correspondence at a granular level to improve the transferability of structural and feature information across domains.
For GCN and GIN, we use the Pytorch Geometric  [^4] https://www.pyg.org/ to implement the model. For other baseline methods, we use the source codes provided by the corresponding paper. For all baseline methods, we vary the dropout rate in the range of {0.1,0.3,0.5,0.7} and then choose the best one. The hidden dimension in these methods is set to 256 for a fair comparison.
Table 7: Statistics of the SEED dataset.
Datasets Graphs Avg. Nodes Avg. Edges Classes
SEED 7,636 62.0 168.2 3

### v4

@@ Appendix E Generalization Bound with Pseudo-label Distillation Module
where (a) is the outcome of applying the union bound with coefficient $\frac{N^{\prime}_{T}}{N_{S}+N^{\prime}_{T}}$, $\frac{N_{S}}{N_{S}+N^{\prime}_{T}}$ respectively; (b) additionally adopt the assumption $N^{\prime}_{T}\ll N_{S}$, following the sleight-of-hand in [Li et al., 2021] Theorem 3.2.
@@
$+C\sqrt{\frac{ln(2/\delta)}{N_{S}}}+2C_{f}C_{g}W_{1}(\mathbb{P}_{S}(G),\mathbb{P}_{T}(G))+\omega,$
where $\omega=\min\left(|\epsilon_{S}(h,\hat{h}_{S}(\mathbf{X})))-\epsilon_{S}(h,\hat{h}_{T}(\mathbf{X})))|,|\epsilon_{T}(h,\hat{h}_{S}(\mathbf{X})))-\epsilon_{T}(h,\hat{h}_{T}(\mathbf{X})))|\right)$, $\hat{\epsilon}_{T}$ is the empirical risk on the high confidence samples, $\epsilon_{T}$ is the empirical risk on the target domain. Besides, we have:
@@
$+2C_{f}C_{g}W_{1}\left(\mathbb{P}_{S}(G),\mathbb{P}_{T}(G)\right)+\omega.$
Appendix F Algorithm
Algorithm 1 Learning Algorithm of DeSGraDA
@@ Appendix F Algorithm
Output: The parameters $\theta$ of Degree-Conscious spiking encoder, parameters $\gamma$ of domain discriminator, and parameter $\eta$ of semantic classifier.
1:Initialize model parameters $\theta$, $\gamma$, and $\eta$
@@
4:     Forward propagate $\mathcal{B}^{s}$ and $\mathcal{B}^{t}$ through the Degree-Conscious spiking encoder
@@
7:     Update model parameters $\theta$, $\gamma$, and $\eta$ via backpropagation
@@
Appendix G Complexity Analysis
Here we analyze the computational complexity of the proposed DeSGraDA. The computational complexity primarily relies on Degree-Conscious spiking representations. For a given graph $G$, $\|A\|_{0}$ denotes the number of nonzeros in the adjacency matrix of $G$. $d$ is the feature dimension. $L$ denote the layer number of spiking encoder. $|V|$ is the number of nodes. $T$ denotes the number of latency step. The spiking encoder takes $\mathcal{O}\left(T\cdot L\cdot\left(\|A\|_{0}\cdot d+|V|\cdot d^{2}\right)\right)$ computational time for each graph. As a result, the complexity of our DeSGraDA is proportional to both $|V|$ and $\|A\|_{0}$.
Figure 5: Visualization of different distributions on the PROTEINS dataset.
Figure 6: Visualization of different distributions on the DD dataset.
Figure 7: Visualization of edge density distributions on the BCI (left) and EEG (right) datasets.
Appendix H Dataset
@@
SEED 3,818 62.00 125.74 3
BCI 1,440 22.00 119.80 2
PROTEINS 1,113 39.10 72.80 2
DD 1,178 284.32 715.66 2
COX2 467 41.22 43.45 2
COX2_MD 303 26.28 335.12 2
BZR 405 35.75 38.36 2
BZR_MD 306 21.30 225.06 2
H.1 Dataset Description
@@
We conduct extensive experiments on different types of datasets. The dataset statistics can be found in Table 4, and their details are shown as follows:
  • SEED. The SJTU Emotion EEG Dataset (SEED) [Zheng and Lu, 2015, Duan et al., 2013] is a widely used Electroencephalography (EEG) benchmark for emotion recognition. It contains EEG recordings from 15 participants watching emotional movie clips that evoke positive, neutral, and negative emotions. Each recording consists of 62-channel EEG signals sampled at 1000 Hz. Based on the edge density, we partition the SEED dataset into three sub-datasets, namely S0, S1, and S2. The sub-datasets exhibit substantial domain disparities among them.
  • BCI. The Brain-Computer Interface (BCI) Competition IV-2a dataset Brunner et al. [2008] is a widely used benchmark for motor imagery EEG classification. It includes EEG recordings from 9 subjects performing four motor imagery tasks: left hand, right hand, feet, and tongue. Each subject completed two sessions on different days, with signals recorded from 22 EEG channels at a sampling rate of 250 Hz. Based on the edge density, we partition the dataset into three parts: B0, B1, and B2.
  • PROTEINS and DD. The PROTEINS [Dobson and Doig, 2003] and DD datasets comprise protein structure graphs commonly used for graph classification tasks. In both datasets, nodes represent amino acids, forming edges between spatially or sequentially adjacent residues. In PROTEINS, each graph is labeled to indicate whether the protein is an enzyme, with edges defined between residues less than 6 Angstroms apart. The DD dataset, derived from the Protein Data Bank, focuses on classifying proteins by structural class and typically exhibits denser graph connectivity than PROTEINS. Additionally, we partition the PROTEINS and DD datasets into four parts based on edge density and node density, P0 to P3 and D0 to D3, respectively.
  • COX2 and COX2_MD. The COX2 dataset Sutherland et al. [2003] consists of 467 molecular graphs, while COX2_MD contains 303 structurally modified counterparts. In both datasets, nodes represent atoms and edges correspond to chemical bonds. Specifically, COX2_MD introduces structural variations to the original COX2 molecules while preserving their semantic labels.
  • BZR and BZR_MD. The BZR dataset Sutherland et al. [2003] comprises 405 molecular graphs, while BZR_MD includes 306 structurally modified graphs derived from BZR. In both datasets, nodes represent atoms and edges denote chemical bonds. BZR_MD introduces structural variations to simulate domain shifts while preserving the original label semantics.
H.2 Data processing
In our implementation, we process the above datasets as follows:
  • For datasets from TUDataset  [^3] https://chrsmrrs.github.io/datasets/, including PROTEINS, DD, BZR, BZR_MD, COX2, and COX2_MD, we utilize the TUDataset module from PyTorch Geometric for loading. Self-loops are added during the preprocessing stage to prevent isolated nodes.
  • For the SEED dataset, we utilize the TorchEEG library [^4] https://torcheeg.readthedocs.io/en/latest/ to transform raw EEG signals into graph-structured data. During graph construction, edges are removed following the approach described in Klepl et al. [2022].
  • For the BCI dataset, we focus on a binary classification subset involving left-hand and right-hand motor imagery tasks, a widely adopted evaluation setting in BCI research. Following the construction protocols proposed in Altaheri et al. [2022, 2023], we randomly remove a portion of edges from each graph during preprocessing.
Appendix I Baselines
Graph kernel method. We compare DeSGraDA with one graph kernel method:
  • WL subtree: Weisfeiler-Lehman (WL) subtree [Shervashidze et al., 2011] is a graph kernel method, which calculates the graph similarity by a kernel function, where it encodes local neighborhood structures into subtree patterns, efficiently capturing the topology information contained in graphs.
Graph-based methods. We compare DeSGraDA with four widely used graph-based methods:
  • GCN: GCN [Kipf and Welling, 2017] is a spectral-based neural network that iteratively updates node representations by aggregating information from neighboring nodes, effectively capturing both local graph structure and node features.
  • GIN: GIN [Xu et al., 2018] is a message-passing neural network designed to distinguish graph structures using an injective aggregation function, theoretically achieving the expressive power of the Weisfeiler-Lehman test.
  • CIN: CIN [Bodnar et al., 2021] extends the Weisfeiler-Lehman framework by integrating cellular complexes into graph neural networks, allowing for the capture of higher-dimensional topological features.
  • GMT: GMT [Baek et al., 2021] utilizes self-attention mechanisms to dynamically adjust the importance of nodes based on their structural dependencies, thereby enhancing both adaptability and performance.
Spiking-based graph methods. We compare DeSGraDA with two spiking-based graph methods:
  • SpikeGCN: SpikeGCN [Zhu et al., 2022] introduces an end-to-end framework designed to integrate the fidelity characteristics of SNNs with graph node representations.
  • DRSGNN: DRSGNN [Zhao et al., 2024] dynamically adapts to evolving graph structures and relationships through a novel architecture that updates node representations in real-time.
Domain adaption methods. We compare DeSGraDA with two recent domain adaption methods:
  • CDAN: CDAN [Long et al., 2018] employs a conditional adversarial learning strategy to reduce domain discrepancy by conditioning adversarial adaptation on discriminative information from multiple domains.
  • ToAlign: ToAlign [Wei et al., 2021b] uses token-level alignment strategies within Transformer architectures to enhance cross-lingual transfer, optimizing the alignment of semantic representations across languages.
  • MetaAlign: MetaAlign [Wei et al., 2021a] is a meta-learning framework for domain adaptation that dynamically aligns feature distributions across domains by learning domain-invariant representations.
Graph domain adaptation methods. We compare DeSGraDA with six graph domain adaption methods:
  • DEAL: DEAL [Yin et al., 2022] uses domain adversarial learning to align graph representations across different domains without labeled data, overcoming discrepancies between the source and target domains.
  • CoCo: CoCo [Yin et al., 2023] leverages contrastive learning to align graph representations between source and target domains, enhancing domain adaptation by promoting intra-domain cohesion and inter-domain separation in an unsupervised manner.
  • SGDA: SGDA [Qiao et al., 2023] utilizes labeled data from the source domain and a limited amount of labeled data from the target domain to learn domain-invariant graph representations.
  • StruRW: StruRW [Liu et al., 2023] introduces a structural re-weighting mechanism that dynamically adjusts the importance of nodes and edges based on their domain relevance. It enhances feature alignment by emphasizing transferable structures while suppressing domain-specific noise.
  • A2GNN: A2GNN [Liu et al., 2024a] introduces a novel propagation mechanism to enhance feature transferability across domains, improving the alignment of graph structures and node features in an unsupervised setting.
  • PA-BOTH: PA-BOTH [Liu et al., 2024b] aligns node pairs between source and target graphs, optimizing feature correspondence at a granular level to improve the transferability of structural and feature information across domains.
DeSGraDA DEAL CoCo SGDA StruRW A2GNN PA-BOTH
PROTEINS 1.0 1.2 1.2 1.0 1.2 22.3 1.7
DD 5.6 6.4 2.5 3.9 4.5 35.1 16.8
SEED 1.1 1.4 1.5 2.6 0.8 16.8 2.8
BCI 0.7 0.8 0.7 0.7 0.7 14.5 1.6
DeSGraDA DEAL CoCo SGDA StruRW A2GNN PA-BOTH
PROTEINS 0.195 0.103 22.123 0.088 0.088 1.313 0.949
DD 0.427 0.400 184.015 0.135 0.140 2.263 0.787
SEED 0.192 0.137 26.187 0.126 0.075 1.414 0.311
BCI 0.224 0.211 35.162 0.103 0.086 0.781 0.123
Appendix J Implementation Details
DeSGraDA and all baseline models are implemented using PyTorch [^5] https://pytorch.org/ and PyTorch Geometric [^6] https://www.pyg.org/. We conduct experiments for DeSGraDA and all baselines on NVIDIA A100 GPUs for a fair comparison, where the learning rate of Adam optimizer set to ${10^{-4}}$, hidden embedding dimension 256, weight decay ${10^{-12}}$, and GNN layers 4. Additionally, DeSGraDA and all baseline models are trained using all labeled source samples and evaluated on unlabeled target samples [Wu et al., 2020]. The performances of all models are measured and averaged on all samples for five different runs.
Appendix K More experimental results
K.1 More performance comparison
In this part, we provide additional results for our proposed method DeSGraDA compared with all baseline models across various datasets, as illustrated in Table 11-14. These results consistently show that DeSGraDA outperforms the baselines in most cases, validating the superiority of our proposed method.
K.2 Training time and memory comparison
We provide detailed comparisons of GPU memory consumption and training time per epoch for DeSGraDA and other graph domain adaptation methods under identical experimental settings in this part, as shown in Tables 5 and 6. It is worth noting that the training phase is typically conducted on more powerful hardware to achieve optimal performance within a reasonable time frame.
Table 7: The results of DeSGraDA with different widely used graph neural networks (GIN, GCN and SAGE) on the PROTEINS and DD dataset. Bold results indicate the best performance.
Methods PROTEINS DD
@@ Appendix I Baselines
DeSGraDA w GCN 71.9 76.8 64.8 68.6 58.2 70.2 57.6 64.1
DeSGraDA w SAGE 73.9 79.8 65.5 74.5 57.8 71.6 59.1 66.7
DeSGraDA w GIN 74.1 81.5 66.4 78.7 59.3 73.0 61.1 68.8
DeSGraDA 76.3 84.6 69.2 83.6 60.1 76.1 63.9 71.7
Table 8: The results of DeSGraDA with different widely used graph neural networks (GIN, GCN and SAGE) on the SEED and BCI dataset. Bold results indicate the best performance.
Methods SEED BCI
@@
DeSGraDA w GCN 55.7 54.5 53.7 54.0 52.8 51.6 52.2 54.2
DeSGraDA w SAGE 55.3 54.5 54.2 54.9 52.7 52.2 52.8 54.7
DeSGraDA w GIN 56.6 57.1 55.8 56.9 53.3 52.9 53.8 55.4
DeSGraDA 58.0 58.2 57.0 58.3 54.1 53.6 54.9 56.2
K.3 More Ablation study
To validate the effectiveness of the different components in DeSGraDA, we conduct more experiments with four variants on DD, SEED, and BCI datasets, i.e., DeSGraDA w CDAN, DeSGraDA w/o PL, DeSGraDA w/o CF and DeSGraDA w/o TL. The results are shown in Table 9 and 10. From the results, we have similar observations as summarized in Section 5.3.
Additionally, we conduct ablation studies to examine the effect of directly replacing the SGNs with commonly used Graph Neural Networks (GNNs) for generating representations for DeSGraDA: (1) DeSGraDA w GCN: It replaces SGNs with GCN [Kipf and Welling, 2017]; (2) DeSGraDA w GIN: It replaces SGNs with GIN [Xu et al., 2018]; (3) DeSGraDA w SAGE: It replaces SGNs with GraphSAGE [Hamilton et al., 2017]. The experimental results across the PROTEINS, DD, SEED, and BCI datasets are shown in Table 7 and 8. However, the critical aspect of our work lies in the specific problem we set up, i.e., low-power and distribution shift environments. In this context, directly replacing SGNs with commonly used GNNs like GIN or GCN is not feasible, as these models are unsuitable for deployment on low-energy devices. As demonstrated in Section 5.4, GNN-based methods have much higher energy consumption than the spike-based methods.
K.4 More Sensitivity Analysis
In this part, we provide additional sensitivity analysis of the proposed DeSGraDA with respect to the impact of its hyperparameters: the latency step $T$ and initial threshold value $V_{th}^{degree}$ in SNNs on the DD, SEED and BCI datasets. The results are illustrated in Figure 8 and 9, where we observe trends similar to those discussed in Section 5.5.
(a) DD
(b) SEED
(c) BCI
Figure 8: Hyperparameter sensitivity analysis of latency step $\mathit{T}$ in SNNs on the DD, SEED and BCI datasets.
(a) DD
(b) SEED
(c) BCI
Figure 9: Hyperparameter sensitivity analysis of the initial threshold $V_{th}^{degree}$ in SNNs on the DD, SEED and BCI datasets.
Table 9: The results of ablation studies on the DD dataset (source → target). Bold results indicate the best performance.
@@
DeSGraDA w CDAN 57.3 72.7 59.9 68.8 57.4 58.1 68.1 60.9 75.9 59.3 80.2 77.9
DeSGraDA w/o PL 58.3 71.8 58.1 69.8 57.9 58.2 67.1 59.3 75.1 60.7 79.1 76.8
DeSGraDA w/o CF 51.0 62.0 57.1 61.8 54.0 51.5 59.0 54.3 62.3 52.9 64.1 69.3
DeSGraDA w/o TL 58.0 72.0 60.4 68.2 58.4 58.8 68.7 60.8 77.6 60.5 76.6 77.4
DeSGraDA 60.1 76.1 63.9 71.7 61.1 62.3 73.6 64.2 80.9 64.7 82.0 80.6
Table 10: The results of ablation studies on the SEED and BCI datasets (source → target). Bold results indicate the best performance.
Methods SEED BCI